1/1

Edgar Yipsel Harburg,

Somewhere over the Rainbow

Somewhere over the rainbow, Way up high There’s a land that I dreamed of Once in a lullaby.

Somewhere over the rainbow, Skies are blue And the dreams that you dare to dream Really do come true.

Some day I’ll wish upon a star And wake up where the clouds are far behind me. Where troubles melt like lemon drops Away above the chimney tops That’s where you’ll find me.

Somewhere over the rainbow, Blue birds fly Birds fly over the rainbow Why, then, oh, why can’t I?

If happy little bluebirds fly Beyond the rainbow Why, oh, why can’t I?

Verkinto de tiu ĉi Angla poemo estas Edgar Yipsel Harburg (civila nomo: Isidore Hochberg, ∗1896-04-08 – †1981-03-05).

Arg-979-1975 (2014-03-14 17:04:48)

La  teksto  troviĝas  interalie  en  http: // artists. letssingit. com/

e-y-h rburg-lyrics-over-the-rainbow-2pp347l# axzz2vx8pOsV7 .

http://www.poezio.net/version?poem-id=979&version-id=,1975